# primordial black holes; microlensing; dark matter

### v1

@@ 3.4 OGLE
Niikura et al. Niikura et al. (2019) used OGLE microlensing observations towards the Galactic bulge to probe planetary mass COs. They found that most events were well-fit by microlensing by stellar populations, and excluded $f>0.1$ for $10^{-6}\lesssim M/M_{\odot}\lesssim 10^{-2}$. There was a distinct population of six ultra-short ($\sim 0.1$ day) events. These events could be due to free-floating planets, however they argued that these events are also consistent with Earth mass PBHs making up a few per-cent of the MW halo. From 2022-2024 OGLE carried out a high-cadence survey of the MC, observing 35 million stars up to 21 times per night Mróz et al. (2024). They only found one event, with Einstein radius crossing time $\sim 90$ days, which allowed them to exclude COs with mass $10^{-8}\lesssim M/M_{\odot}\lesssim 0.01$ making up more than $1\%$ of the MW halo.
3.5 M31
@@
More recently high cadence (2 minute sampling interval) observations of M31 have been carried out using the Hyper Suprime-Cam (HSC) on the Subaru telescope Niikura et al. (2019); Sugiyama et al. (2026). Initial observations found one candidate event which led to a constraint $f\lesssim 10^{-2}$ for $10^{-5}\lesssim M/M_{\odot}\lesssim 10^{-9}$ Niikura et al. (2019, 2019). The finite size of the lensed stars reduces the magnification Witt and Mao (1994), and makes the constraint weaker than it would otherwise be, for $M\lesssim 10^{-7}M_{\odot}$ Niikura et al. (2019); Smyth et al. (2020). Very recently this work has been updated with additional data, and improved analysis of the original data (in particular the inclusion of finite source size effects in the detection efficiency estimation) Sugiyama et al. (2026). Twelve microlensing candidates have been identified, with 4 being classed as secure, high-significance, detections. If these events are due to COs in the MW halo the corresponding mass and halo fraction are $M\sim(10^{-7}-10^{-6})M_{\odot}$ and $f\sim 0.01-0.1$.
@@ 3.7 Summary
Figure 4: Constraints on the CO mass, $M$, and MW halo fraction, $f$, at 95% confidence, assuming all CO have the same mass. The solid lines show the exclusion limits from EROS (’E’) Tisserand et al. (2007) (yellow), MACHO long timescale (’M-long’) Allsman et al. (2001) (red), EROS plus MACHO long timescale (’E+M’) Blaineau et al. (2022) (orange), OGLE Mróz et al. (2024) and OGLE high cadence (’OGLE-hc’) Mróz et al. (2024) (blue), and Kepler (’K’) Griest et al. (2014) (green). The purple lines are from high cadence observations of M31 using the Subaru-HSC. ’HSC19’ shows the original exclusion limit Niikura et al. (2019) as calculated in Ref. Croon et al. (2020), while ’HSC26’ shows the exclusion limit from the latest analysis, under the assumption that none of the events observed are due to CO in the MW halo. The dashed line shows the allowed parameter values assuming that the 4 ’secure’ high-significance events are due to CO in the MW halo. The dashed red line shows the allowed parameter values from the 13 set A events in the final 5.7 year MACHO data set, as published in Ref. Alcock et al. (2000). Subsequent observations indicate that some of these events are not microlensing (see Sec. 3.2 and Refs. Tisserand et al. (2007); Mróz et al. (2025) for discussion). The dotted blue line, labeled ’OGLE?’, shows the parameter values found assuming that the ultra-short duration events observed towards the Galactic bulge by OGLE are due to COs in the MW halo Niikura et al. (2019). The dot-dashed black line shows the envelope of the tightest constraint for each value of $M$. This plot was made using a modified version of Kavanagh’s PBHBounds Kavanagh .
Figure 2 shows the final published parameter constraints on the halo fraction, $f$, and compact object mass, $M$, (assuming all compact objects have the same mass) from the surveys discussed in this section  [^5] Results from M31 surveys, other than HSC, are not shown. These surveys quoted approximate allowed or excluded values for the halo fraction, but they did not publish exclusion limit curves or allowed regions of parameter space.. Table 2 contains a summary of the key features of the MACHO Alcock et al. (2000), EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024) LMC microlensing surveys. See Table 1 of Ref. Moniez (2010) for a more detailed comparison of stellar microlensing surveys (as of 2010).
@@ 4.1 Observations
Following early work in Ref. Cieplak and Griest (2013), Ref. DeRocco et al. (2024) showed that the Roman Space Telescope’s Galactic Bulge Time Domain Survey will be able to to discriminate planetary-mass PBHs from free floating planets, and probe halo fractions $f\sim 10^{-3}-10^{-4}$ using the timescales of (photometric) microlensing events. Ref. Fardeen et al. (2024) showed that this survey could probe $f\sim 10^{-3}$ for $10^{-1}\lesssim M/M_{\odot}\lesssim 10^{2}$ if it is possible to distinguish PBHs and stellar lenses using the joint distributions of the timescales and angular Einstein radii of events. If it is not possible to do this then the sensitivity will be reduced to $f\sim 10^{-1}$, weaker than existing OGLE limits in this mass range Mróz et al. (2024). Similar constraints may also be possible using the Rubin Observatory Drlica-Wagner et al. (2019). With improved understanding of backgrounds, Gaia time-series data could probe halo fractions down to $f\sim 10^{-3}$ for multi-Solar mass PBHs Verma and Rentala (2023), see also Ref. Belokurov and Evans (2002). Thirty-metre class, Extremely Large Telescopes, will also be able to probe PBHs via astrometric microlensing Bird et al. (2023). Combining photometric and astrometric microlensing data would aid the detection and characterization of events Pruett et al. (2024). In all cases, correctly identifying events and minimizing the false positive rate will be crucial (as we have seen for previous surveys in Sec. 3). See Ref. Crispim Romao et al. (2025) for work in this direction.
@@ 4.2 Theory
The key theoretical uncertainty is the distribution of PBHs, specifically their local velocity distribution and also whether PBHs form compact clusters. As discussed in Sec. 2.2, uncertainties in the DM distribution lead to uncertainties in the differential event rate and hence the constraints on the halo fraction. If the DM is smoothly distributed, the most significant uncertainty is the dependence of the rate of short-timescale events, and hence the constraints on $f$ for large masses, on the local velocity distribution Green (2025). If PBHs form compact clusters, such that the cluster as a whole lenses, the constraints would change significantly Calcino et al. (2018). Understanding the present day spatial distribution of PBHs formed from realistic density perturbation distributions (i.e. how many PBHs are in clusters, and what are their properties? how many PBHs are in binaries, and what are their properties?) is a major outstanding challenge.
@@ 5 Summary
We have explored how stellar microlensing surveys probe compact object dark matter, in particular primordial black holes. Early indications from the MACHO collaboration that a significant fraction of the Milky Way halo could be in the form of roughly Solar Mass COs Alcock et al. (1997, 2000) were not confirmed by EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024). Long duration Mróz et al. (2024) and high cadence Niikura et al. (2019); Mróz et al. (2024) observations have extended the range of CO masses probed by microlensing. Under the standard assumptions outlined in Sec. 2, collectively these surveys exclude COs with mass $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$ making up all of the DM in the MW, while those in the mass range $10^{-10}\lesssim M/M_{\odot}\lesssim 1$ can make up less than one percent. Further theoretical work is required to ascertain how robust the standard assumptions are, in particular regarding the spatial distribution of PBHs. Future observations, in particular astrometric ones using the Roman Telescope and Rubin Observatory, can potentially probe PBHs down to $f\sim 10^{-3}-10^{-4}$. However further work is required to develop techniques for discriminating PBH-induced lensing events from backgrounds (including, but not limited to, lensing events due to astrophysical objects).

### v2

@@ 3.4 OGLE
Niikura et al. Niikura et al. (2019) used OGLE microlensing observations towards the Galactic bulge to probe planetary mass COs. They found that most events were well-fit by microlensing by stellar populations, and excluded $f>0.1$ for $10^{-6}\lesssim M/M_{\odot}\lesssim 10^{-2}$. There was a distinct population of six ultra-short ($\sim 0.1$ day) events. These events could be due to free-floating planets, however they argued that these events are also consistent with Earth mass COs making up a few per-cent of the MW halo. From 2022-2024 OGLE carried out a high-cadence survey of the MC, observing 35 million stars up to 21 times per night Mróz et al. (2024). They only found one event, with Einstein radius crossing time $\sim 90$ days, which allowed them to exclude COs with mass $10^{-8}\lesssim M/M_{\odot}\lesssim 0.01$ making up more than $1\%$ of the MW halo.
3.5 Surveys focused on M31
@@
More recently high cadence (2 minute sampling interval) observations of M31 have been carried out using the Hyper Suprime-Cam (HSC) on the Subaru telescope Niikura et al. (2019); Sugiyama et al. (2026). Initial observations found one candidate event which led to a constraint $f\lesssim 10^{-2}$ for $10^{-9}\lesssim M/M_{\odot}\lesssim 10^{-5}$ Niikura et al. (2019, 2019). The finite size of the lensed stars reduces the magnification Witt and Mao (1994), and makes the constraint weaker than it would otherwise be, for $M\lesssim 10^{-7}M_{\odot}$ Niikura et al. (2019); Smyth et al. (2020). Very recently this work has been updated with additional data, and improved analysis of the original data (in particular the inclusion of finite source size effects in the detection efficiency estimation) Sugiyama et al. (2026). Twelve microlensing candidates have been identified, with 4 being classed as secure, high-significance, detections. If these events are due to COs in the MW halo the corresponding mass and halo fraction are $M\sim(10^{-7}-10^{-6})M_{\odot}$ and $f\sim 0.01-0.1$. Mróz and Udalski have, however, carried out an independent reanalysis of the data and argue that all 12 candidate events are in fact variable stars Mróz and Udalski (2026).
@@ 3.7 Summary
Figure 4: Constraints on the CO mass, $M$, and MW halo fraction, $f$, at 95% confidence, assuming all CO have the same mass. The solid lines denote exclusion limits, while the dot-dashed black line shows the envelope of the tightest constraint for each value of $M$. The dashed lines show allowed regions of parameter space. The exclusion limits are from EROS (’E’) Tisserand et al. (2007) (yellow), MACHO long timescale (’M-long’) Allsman et al. (2001) (red), EROS plus MACHO long timescale (’E+M’) Blaineau et al. (2022) (orange), OGLE Mróz et al. (2024) and OGLE high cadence (’OGLE-hc’) Mróz et al. (2024) (blue), and Kepler (’K’) Griest et al. (2014) (green). The purple lines are from high cadence observations of M31 using the Subaru-HSC. ’HSC19’ shows the original exclusion limit Niikura et al. (2019) as calculated in Ref. Croon et al. (2020), while ’HSC26’ shows the exclusion limit from the latest analysis, under the assumption that none of the events observed are due to CO in the MW halo. The dashed line shows the allowed parameter values assuming that the 4 ’secure’ high-significance events are due to CO in the MW halo. The dashed red line shows the allowed parameter values from the 13 set A events in the final 5.7 year MACHO data set, as published in Ref. Alcock et al. (2000). Subsequent observations indicate that some of these events are not microlensing (see Sec. 3.2 and Refs. Tisserand et al. (2007); Mróz et al. (2025) for discussion). The dotted blue line, labeled ’OGLE?’, shows the parameter values found assuming that the ultra-short duration events observed towards the Galactic bulge by OGLE are due to COs in the MW halo Niikura et al. (2019). This plot was made using a modified version of Kavanagh’s PBHBounds Kavanagh .
Figure 4 shows the final published parameter constraints on the halo fraction, $f$, and compact object mass, $M$, (assuming all compact objects have the same mass) from the surveys discussed in this section  [^7] Results from M31 surveys, other than HSC, are not shown. These surveys quoted approximate allowed or excluded values for the halo fraction, but they did not publish exclusion limit curves or allowed regions of parameter space.. The solid lines denote exclusion limits, while the dot-dashed black line shows the envelope of the tightest constraint for each value of $M$. The dashed lines show allowed regions of parameter space. Table 2 contains a summary of the key features of the MACHO Alcock et al. (2000), EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024) LMC microlensing surveys. See Table 1 of Ref. Moniez (2010) for a more detailed comparison of stellar microlensing surveys (as of 2010).
@@ 4.1 Observations
To probe CO DM with the Roman Space Telescope’s Galactic Bulge Time Domain Survey it will be necessary to discriminate events due to CO in the MW halo from events due to astrophysical objects in the luminous components of the MW. Following early work in Ref. Cieplak and Griest (2013), Ref. DeRocco et al. (2024) showed that it will be possible to discriminate planetary-mass CO DM from free-floating planets, and probe halo fractions $f\sim 10^{-3}-10^{-4}$ using just the timescales of (photometric) microlensing events. Ref. Fardeen et al. (2024) showed that this survey could probe $f\sim 10^{-3}$ for $10^{-1}\lesssim M/M_{\odot}\lesssim 10^{2}$ if it is possible to distinguish CO DM and stellar lenses using the joint distributions of the timescales and angular Einstein radii of events. If it is not possible to do this then the sensitivity will be reduced to $f\sim 10^{-1}$, weaker than existing OGLE limits in this mass range Mróz et al. (2024). Similar constraints may also be possible using the Rubin Observatory Drlica-Wagner et al. (2019). With improved understanding of backgrounds, Gaia time-series data could probe halo fractions down to $f\sim 10^{-3}$ for multi-Solar mass CO DM Verma and Rentala (2023), see also Ref. Belokurov and Evans (2002). Thirty-metre class, Extremely Large Telescopes, will also be able to probe CO DM via astrometric microlensing Bird et al. (2023). Combining photometric and astrometric microlensing data would aid the detection and characterization of events Pruett et al. (2024). In all cases, correctly identifying events and minimizing the false positive rate will be crucial (as we have seen for previous surveys in Sec. 3). See Ref. Crispim Romao et al. (2025) for work in this direction.
@@ 4.2 Theory
As discussed in Sec. 2.2, uncertainties in the DM distribution lead to uncertainties in the differential event rate and hence the constraints on the halo fraction. If the DM is smoothly distributed, the most significant uncertainty is the dependence of the rate of short-timescale events, and hence the constraints on $f$ for large masses, on the local velocity distribution Green (2025). In the case of PBHs, if a significant fraction are in compact clusters, such that the cluster as a whole lenses, the constraints would change significantly Calcino et al. (2018). Understanding the present day spatial distribution of PBHs formed from realistic density perturbation distributions (i.e. how many PBHs are in clusters, and what are their properties? how many PBHs are in binaries, and what are their properties?) is a major outstanding challenge.
@@ 5 Summary
We have explored how stellar microlensing surveys probe compact object dark matter, in particular primordial black holes. Early indications from the MACHO collaboration that a significant fraction of the Milky Way halo could be in the form of roughly Solar Mass COs Alcock et al. (1997, 2000) were not confirmed by EROS Tisserand et al. (2007) and OGLE Mróz et al. (2024). Long duration Mróz et al. (2024) and high cadence Niikura et al. (2019); Mróz et al. (2024) observations have extended the range of CO masses probed by microlensing. Under the standard assumptions outlined in Sec. 2, collectively these surveys exclude COs with mass $10^{-11}\lesssim M/M_{\odot}\lesssim 10^{4}$ making up all of the DM in the MW, while those in the mass range $10^{-10}\lesssim M/M_{\odot}\lesssim 1$ can make up less than one percent. Further theoretical work is required to ascertain how robust the standard assumptions regarding the CO distribution are. For PBHs a key question is whether a significant fraction are in compact clusters. Future observations, in particular astrometric ones using the Roman Telescope and Rubin Observatory, can potentially probe CO dark matter down to $f\sim 10^{-3}-10^{-4}$. However further work is required to develop techniques for discriminating CO dark matter-induced lensing events from backgrounds, including lensing events due to known astrophysical objects, such as stars and planets.